### v1

# Large Language Models for Scholarly Ontology Generation: An Extensive Analysis in the Engineering Field

Abstract
Ontologies of research topics are crucial for structuring scientific knowledge, enabling scientists to navigate vast amounts of research, and forming the backbone of intelligent systems such as search engines and recommendation systems. However, manual creation of these ontologies is expensive, slow, and often results in outdated and overly general representations. As a solution, researchers have been investigating ways to automate or semi-automate the process of generating these ontologies. This paper offers a comprehensive analysis of the ability of large language models (LLMs) to identify semantic relationships between different research topics, which is a critical step in the development of such ontologies. To this end, we developed a gold standard based on the IEEE Thesaurus to evaluate the task of identifying four types of relationships between pairs of topics: broader, narrower, same-as, and other. Our study evaluates the performance of seventeen LLMs, which differ in scale, accessibility (open vs. proprietary), and model type (full vs. quantised), while also assessing four zero-shot reasoning strategies. Several models have achieved outstanding results, including Mixtral-8×7B, Dolphin-Mistral-7B, and Claude 3 Sonnet, with F1-scores of 0.847, 0.920, and 0.967, respectively. Furthermore, our findings demonstrate that smaller, quantised models, when optimised through prompt engineering, can deliver performance comparable to much larger proprietary models, while requiring significantly fewer computational resources.
keywords:
Large Language Models , Zero-Shot Learning , Ontology Generation , Research Topics , Scholarly Knowledge , Scientific Knowledge Graphs1 Introduction
Ontologies of research topics, along with other knowledge organization systems (KOS) such as taxonomies and thesauri, are essential for organising and retrieving information from digital libraries. Indeed, major publishers such as Elsevier, SpringerNature, ACM, IEEE, and PubMed, utilise a range of organization systems like the All Science Journal Classification (ASJC), SpringerNature Taxonomy, ACM Computing Classification System111ACM Computing Classification System - https://dl.acm.org/ccs, the IEEE Theasurus, and Medical Subject Headings, to organise their published content [1, 2, 3] and ensuring efficient scientific dissemination. Furthermore, ontologies of research topics play a vital role in supporting a wide range of intelligent systems to navigate and interpret academic literature [4, 5, 6]. These include search engines [6], conversational agents [7], analytics dashboards [8], and recommender systems [5].
Nevertheless, generating and maintaining ontologies of research topics is both expensive and time-consuming, typically requiring the involvement of several domain experts over an extended period of time [9]. Moreover, the ever-increasing volume of literature, with approximately 2.5 million new scientific papers published annually [10, 11], necessitates more frequent updates to ensure these ontologies accurately reflect the latest research developments. Over the past 25 years, Artificial Intelligence (AI) researchers have concentrated on developing methods for the automated or semi-automated generation of such ontologies, with varying degrees of success [12, 13, 9, 14]. Some notable achievements include the semi-automatic pipeline for generating OpenAlex Topics [15] and the automatic development of the Computer Science Ontology [16], along with a few other initiatives. However, most of those ontologies are still manually created because current systems struggle to generate large-scale, detailed representations of research topics. Consequently, this challenge remains far from being fully addressed.
Over the past two years, Large Language Models (LLMs) have transformed the field of Natural Language Processing (NLP) by significantly improving machine comprehension of human language. In particular, some studies have demonstrated that LLMs can achieve remarkable performance even in zero-shot settings, where the model makes predictions without prior exposure to specific tasks or examples [17].
In this paper, we investigate whether LLMs can effectively address the challenges involved in generating ontologies of research topics. Specifically, we conduct a comprehensive analysis of the ability of modern LLMs to identify semantic relationships between research topics. To this end, we introduce IEEE-Rel-1K, a gold standard derived from the IEEE Thesaurus, designed to evaluate the identification of four types of relationships (broader, narrower, same-as, and other) between 1,000 pairs of topics. We evaluated 17 off-the-shelf LLMs on this dataset, varying in scale, accessibility (open or proprietary), and model type (full or quantised). The models included 3 full open models, 10 open quantised models, and 4 proprietary models. For each model, we employed four distinct zero-shot prompting strategies. The first strategy involved a straightforward prompt asking the model to determine the relationship between two topics, and . The second strategy extended this by applying the same method to both the original pair (, ) and the reversed pair (, ), with the results then being combined. The third strategy utilised a chain-of-thought prompt, encouraging the model to reason through the relationship step by step. The fourth strategy was a two-way version of this chain-of-thought approach, where the process is repeated for both the original and reversed pairs before integrating the outcomes.
Our findings indicate that several models exhibit exceptional zero-shot reasoning capabilities for this task. Proprietary models led the way, with Claude 3 Sonnet achieving an impressive 0.967 F1. However, several open models also delivered highly competitive results, particularly when chain-of-thought prompting techniques were employed. Among the open models, the top performer was Dolphin-Mistral-7B (0.920 F1), a Mistral variant trained on the Dolphin dataset. The Mixtral model also exhibited strong performance (0.847 F1). The findings also indicate that prompting strategies can significantly influence model performance. Different prompting strategies led to dramatic variations in performance, with gains of over 0.354 F1 points in some instances.
In summary, the contributions of this paper are as follows:
We conduct a comprehensive analysis of 17 recent LLMs in identifying semantic relationships between research topics in a zero-shot setting.
We introduce IEEE-Rel-1K, a gold standard for research topic relation prediction, designed to facilitate the evaluation of models in this domain.
We explore this task under various conditions, examining different types of LLMs and evaluating the effectiveness of multiple prompting strategies.
We provide the complete codebase for our methodologies, including the gold standard222Gold Standard and code for experiments can be found at https://github.com/ImTanay/LLM-Automatic-Ontology-Generation.
The remainder of this paper is structured as follows. Section 2 provides an overview of the literature. Section 3 defines the task, describes the gold standard, and presents the seventeen LLMs employed in this study. Section 4 describes the experiments and provides details on the implementation setup. Section 5 presents and discusses our results. Section 6 provides an analysis of the types of errors made by the models. Finally, in Section 7 we conclude the paper and provide future directions.
2 Related Work
In this section, we review the literature, by beginning with a discussion on the ontologies of research areas (Section 2.1) and then explore the state-of-the-art methodologies for generating these ontologies (Section 2.2). In this context, an ontology is a formal representation of knowledge within a specific domain, defining the types of entities that exist and the relationships between them. An ontology of research areas applies this concept to the scholarly domain, systematically organising the various fields of study, their sub-disciplines, and their interconnections [18].
2.1 Ontologies of Research Areas
In the literature we can find several ontologies of research areas, each differing in scope, size, depth, and curation. Notable examples include the ACM Computing Classification System, Medical Subject Headings, IEEE Thesaurus, and UNESCO Thesaurus [2, 3, 19, 15].
One crucial aspect of these ontologies is their coverage. For instance, the Medical Subject Headings (MeSH) is a single-field ontology with over 30K concepts in the field of Medicine [2]. It is maintained by the National Library of Medicine (US), and mostly employed to organised content in the Medline digital library. It receives yearly updates with minor amendments to reflect the latest developments of the field.
The field of Computer Science is represented by the ACM Computing Classification System333The ACM Computing Classification System – http://www.acm.org/publications/class-2012 [3] and the Computer Science Ontology (CSO) [19]. The former contains about 2K research topics and it is maintained by the Association for Computing Machinery (ACM) for its digital library. CSO, on the other hand, offers a more comprehensive view with over 14,000 research concepts. CSO was automatically generated using Klink-2 [9], starting from a corpus of 16 million scientific publications; and is being employed by Springer Nature to enhance the metadata quality of their computer science proceedings, including the Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence [20].
In the field of Engineering, there is the IEEE Thesaurus, which contains 12K terms. Curated by the Institute of Electrical and Electronics Engineers (IEEE), it is updated annually and serves as a basis for classifying academic articles and research within the IEEE digital library.
With regard to ontologies that cover multiple fields we can find the UNESCO Thesaurus. Curated by UNESCO, it covers 4.4K subjects and it is mainly employed for indexing and searching resources within the UNESCO’s document repository. It undergoes frequent minor revisions, typically every 2-3 months. Similarly, the ANZSRC Fields of Research (FoR) developed by the Australia and New Zealand research councils, covers around 4.4K topics across a wide range of subjects, but it receives less frequent updates (10 years). It is employed by Digital Science to organise research content within their main datasets: Dimensions.ai and Figshare.
Lastly, OpenAlex Topics is another multi-field taxonomy with 4,800 topics, primarily used within OpenAlex, a free and open catalogue of scholarly articles; and currently at its first release [15].
While most existing ontologies and taxonomies are manually curated, a labour-intensive and costly process, recent automatic and semi-automatic generation methods have led to the creation of CSO and OpenAlex Topics. However, these methods have yet to fully harness the potential of advanced LLMs.
2.2 Generation of Ontologies of research topics
Previous decades have seen the emergence of several automated solutions for ontology creation, yet these have primarily been applied outside the scholarly domain, explaining why scholarly ontologies often remain manually curated. This section will first explore successful general approaches to automatic ontology generation, followed by a look at the limited applications within the scholarly domain.
Early work focused on semi-automatic methods, combining human expertise with computational techniques. Maedche and Staab [21] introduced a method using linguistic and statistical techniques to extract concepts and relationships from text, but heavily relied on human input. Cimiano et al. [22] advanced this with Text2Onto, integrating techniques from machine learning and natural language processing to automate extraction, though it still required substantial human validation. Fully automatic methods emerged with OntoLearn [23], which used lexical resources and statistical methods to reduce human oversight. The same team [24] later improved it with a graph-based algorithm in OntoLearn Reloaded, enhancing accuracy and robustness in complex domains. The integration of deep learning marked a significant advancement for the field of ontology generation. BERT [25] revolutionised language understanding, enhancing concept extraction and relationship identification. BERTopic [26] leveraged embeddings for coherent topic generation, influencing ontology and taxonomy extraction. Le et al. [27] use hyperbolic embeddings to capture the hierarchical structure of concepts. As recent advancements, Chen et al.[28] employed pre-trained models for developing hierarchical taxonomies, obtaining improved results compared to state-of-the-art solutions.
In the last years, we saw the emergence of several approaches specifically tailored for generating ontologies of research topics. For instance, Shen et al. [29] developed a method to discover concepts and arrange them hierarchically. Their resulting extensive taxonomy was then implemented in Microsoft Academic Graph. Another notable method is Klink-2 [9], which extracts large-scale ontologies of research topics and is currently used to generate CSO. The algorithm starts with a set of keywords and analyses their relationship based on the extent of their co-occurrence. It infers semantic relationships between keyword pairs by utilising a modified subsumption method combined with temporal information to assess hierarchical relationships, and topic similarity strategies to identify keywords referring to the same subject. Finally, the OpenAlex team expanded the ASJC structure found in Scopus by incorporating over 4.5K new topics identified through citation clustering [15]. This involved clustering citation networks to identify coherent and thematically related groups of papers, using LLMs to label these groups, and then connecting the newly created concepts to Scopus’ ASJC.
Despite significant advancements in automating ontology generation, the creation of comprehensive and fine-grained ontologies continues to pose a considerable challenge. We hypothesise, however, that the advanced language understanding capabilities of LLMs could play a crucial role in addressing this gap. This paper represents an initial step towards exploring this potential.
3 Background
This section defines the task under investigation (Section 3.1), describes the gold standard developed for this study (Section 3.2), and presents an overview of the LLMs used in the experiments (Section 3.3).
3.1 Task definition
This paper addresses the challenge of identifying semantic relationships between pairs of research topics, denoted in the following as and . We formalise this task as a single-label, multi-class classification problem, where each topic pair is assigned to one of four distinct categories commonly utilised in research topic ontologies [30]: broader, narrower, same-as, or other.
We define the four categories as follows:
broader: is a parent topic of . E.g., machine learning is a broader area than deep neural networks;
narrower: is a child topic of . E.g., distributed databases is a specific area within databases. This is the inverse relationship of broader;
same-as: and can be used interchangeably to refer to same concept. E.g., haptic interface and haptic device. This relationship is symmetric;
other: and do not relate according to the above categories. E.g., software algorithms and cyclones. Also this relationship is symmetric.
3.2 Gold Standard
To evaluate the performance of LLMs on this task, we created a gold standard by selecting a sample of 1,000 semantic relationships from the IEEE Thesaurus. We have named this dataset IEEE-Rel-1K (IEEE Relations). The IEEE thesaurus contains around 11,730 engineering, technical, and scientific terms, including IEEE-specific society terms. It is compiled from IEEE transactions, journal articles, conference papers, and standards, and reflects the vocabulary used within the engineering and scientific fields of IEEE. Specifically, we utilised the IEEE Thesaurus v1.02444A copy of version 1.02 of the IEEE Thesaurus is available at https://github.com/angelosalatino/ieee-taxonomy-thesaurus-rdf/blob/main/source/ieee-thesaurus_2023.pdf., which is dated back to July 2023, and is available as a PDF following the ANSI/NISO Z39.4-2021 standard [31].
To create IEEE-Rel-1K, we developed a Python script555Script for extracting semantic relationships in IEEE Thesaurus - https://github.com/angelosalatino/ieee-taxonomy-thesaurus-rdf. for extracting the hierarchical structure and relationships between terms from the IEEE Thesaurus PDF file. Next, we randomly selected 250 relationships for each of the categories: broader, narrower, and same-as. For same-as, we used a combination of ‘use preferred term’ and ‘used for’ relationships as defined in the ANSI/ISO standard. To complete the dataset, we randomly generated 250 pairs of topics, ensuring they were not semantically related within the IEEE Thesaurus, and labelled these as other.
We released IEEE-Rel-1K within the GitHub repository: https://github.com/ImTanay/LLM-Automatic-Ontology-Generation.
3.3 Large Language Models
Our experiments included a diverse set of seventeen LLMs, all transformer-based and utilising a decoder-only architecture. However, these models exhibit variations in training parameters, quantisation levels, openness, and fine-tuning methodologies. We categorised these models into two primary groups: open (13) and proprietary (4). This distinction was necessary because we lack comprehensive information about proprietary models in terms of their internal structures and any supplementary systems that may support them. As a result, a direct comparison between open and proprietary models would not be entirely fair.
We further divided the 13 open models into two groups: 3 full models and 10 quantised models, the latter being quantised to 8 bits. The quantised models were employed to assess the performance of more computationally efficient solutions. Since some quantised models have been fine-tuned on specific datasets, they may outperform the original models from which they were derived in certain tasks.
Detailed descriptions of each model are provided below.
3.3.1 Open Models
Mistral-7B (shortened as mistral) is trained on Open-Web data, has 7 billion parameters, and utilises grouped-query attention and sliding window attention with a context window of 32,000 tokens.
Mixtral-8×7B (shortened as mixtral) is a sparse mixture-of-experts network incorporating 46.7 billion parameters. Only 12.9 billion parameters are used per token, whereas the remaining parameters are part of the “expert” layers within its sparse mixture-of-experts architecture. It has a context window of 32,000 tokens.
Llama-2-70B (shortened as llama-2) has 70 billion parameters and a context window of 4,096 tokens. It is trained on a corpus that includes 2 trillion tokens of data from publicly available sources.
3.3.2 Quantised Models
Dolphin-2.1-Mistral-7B (shortened as dolphin-mistral666Dolphin-2.1-Mistral-7B - https://huggingface.co/TheBloke/dolphin-2.1-mistral-7B-GGUF) is a model with 7 billion parameters and a token context capacity of 4,096. It is based on Mistral-7B and fine-tuned with the Dolphin777Dolphin Dataset - https://huggingface.co/datasets/cognitivecomputations/dolphin dataset, which is an open-source implementation of Microsoft’s Orca [32], with an addition of Airoboros888Airoboros - https://huggingface.co/datasets/jondurbin/airoboros-2.1 dataset.
Dolphin-2.6-Mistral-7B-dpo-laser (shortened as dolphin-mistral-dpo999Dolphin-2.6-Mistral-7B-dpo-laser - https://huggingface.co/TheBloke/dolphin-2.6-mistral-7B-dpo-laser-GGUF) is based on Mistal-7B with 7 billion parameters and fine-tuned on top of Dolphin DPO using Layer Selective Rank Reduction (LASER) [33]. It offers a context window of 4,096 tokens.
Dolphin2.1-OpenOrca-7B (shortened as dolphin-openorca101010 Dolphin2.1-OpenOrca-7B - https://huggingface.co/TheBloke/Dolphin2.1-OpenOrca-7B-GGUF) is a model that blends Dolphin-2.1-Mistral-7B and Mistral-7B-OpenOrca models. These models were merged using the “ties merge” [34] technique, keeping the same number of training parameters and token context window size, respectively 7 billion and 4,096.
OpenChat-3.5-1210 (shortened as openchat111111OpenChat-3.5-1210 - https://huggingface.co/TheBloke/openchat-3.5-1210-GGUF) is a model fine-tuned on top of Mistral-7B-v0.1. It possesses similar properties and consists of 7 billion parameters with a context window size of 8,192 tokens.
OpenChat-3.5-0106-Gemma (shortened as Openchat-gemma121212OpenChat-3.5-0106-Gemma - https://huggingface.co/gguf/openchat-3.5-0106-gemma-GGUF) is a model trained on openchat-3.5-0106131313openchat-3.5-0106 - https://huggingface.co/openchat/openchat-3.5-0106 data using Conditioned Reinforcement Learning Fine-Tuning (C-RLFT) framework [35, 36]. This model shares the same properties of the openchat-3.5-0106 model, which was fine-tuned141414https://huggingface.co/openchat/openchat-3.5-1210/discussions/4#658288f1168803bdee13d6b3 on top of Mistral-7B. It consists of 7 billion parameters with a context window of 8,192 tokens.
SOLAR-10.7B-Instruct-v1.0 (shortened as solar151515SOLAR-10.7B-Instruct-v1.0 - https://huggingface.co/TheBloke/SOLAR-10.7B-Instruct-v1.0-GGUF) a 10.7 billion parameters model with a Depth Up-Scaling (DUS) architecture that includes architectural modifications and continued pretraining [37]. It is trained on the OpenOrca161616OpenOrca dataset - https://huggingface.co/datasets/Open-Orca/OpenOrca dataset.
Mistral-7B-OpenOrca (shortened as mistral-openorca171717Mistral-7B-OpenOrca - https://huggingface.co/Open-Orca/Mistral-7B-OpenOrca) is based on Mistral-7B-Instruct-v0.1 and has been fine-tuned using the OpenOrca dataset [32]. This dataset is a comprehensive collection of augmented Fine-tuned LAnguage Net (FLAN) data aligned with the distributions outlined in the Orca paper [38, 39]. The model has 7 billion parameters and a context window of 4,096.
Eurus-7b-sft (shortened as eurus181818Eurus-7b-sft - https://huggingface.co/mradermacher/Eurus-7b-sft-GGUF) is fine-tuned on top of Mistral-7B-Instruct-v0.2 using all correct actions in UltraInteract, mixing a small proportion of UltraChat, ShareGPT, and OpenOrca examples, and posses 7 billion parameters with a context window size of 4,096 [40].
Llama-3-8B-Instruct (shortened as llama-3191919Llama-3-8B-Instruct - https://huggingface.co/lmstudio-community/Meta-Llama-3-8B-Instruct-GGUF) has 8 billion parameters and a context length of almost 8,192 tokens [41]. It utilises supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF).
Orca-2-13B (shortened as orca-2202020Orca-2-13B - https://huggingface.co/TheBloke/Orca-2-13B-GGUF) is a model with 13 billion parameters and a 2,048-token context window. It was trained on a mix of publicly available and curated datasets to enhance its reasoning and comprehension abilities. The fine-tuning process was based on datasets like Super-NaturalInstructions and FLAN to ensure strong performance across NLP tasks [42].
3.3.3 Proprietary Models
We considered four proprietary models. Due to undisclosed details, we can only provide limited information. For example, their number of parameters is not officially disclosed.
GPT-3.5-Turbo-Instruct (shortened as gpt-3.5) is trained using RLHF, following similar methods as InstructGPT, but with some differences in the data collection setup. It has a context window size of 4,096 tokens and was trained with data up to September 2021.
GPT-4-Turbo (shortened as gpt-4) has been pre-trained to predict the next token in a document. It was trained using publicly available data and data licensed from third-party providers up to December 2023. The model was then fine-tuned using RLHF. It has a context window of 128,000 tokens.
Claude 3 Haiku (shortened as haiku) is a model with a context size of 200,000 tokens, offering a maximum output of 4,096 tokens with training data up to August 2023.
Claude 3 Sonnet (shortened as sonnet) offers a context size of 200,000 tokens. It has the capability to generate a maximum output of 4,096 tokens and has been trained on data up to August 2023. While haiku is optimised for speed and cost-effectiveness in applications requiring immediate responses, sonnet features a larger model architecture giving it advantage in handling complex tasks that require in-depth understanding of language [43].
4 Experiments
We evaluated the 17 models described in Section 3.3 against IEEE-Rel-1K, the gold standard defined in Section 3.2, employing four different experiments that varied based on the prompting strategy utilised. Figure 1 illustrates the flowcharts of these experiments.
The experiments were categorised into two primary groups: standard prompting and chain-of-thought (CoT) prompting [44]. For each category, we conducted two distinct experiments. The first approach, labelled one-way, involves the model inferring the relationship between and directly. The second approach, referred to as two-way, builds upon the first by having the model initially infer the relationship between and , then infer the relationship between and , and finally combine the results using heuristic rules.
In all cases, the performance of the model was assessed using standard precision, recall and F1.
To summarise, the four experiments can be labelled as follows:
In the following, we will describe them in detail.
4.1 Standard prompting
In the experiments using standard prompting (Fig. 1(a)), we generated a prompt for each pair of research topics from the gold standard using a predefined template. The prompt template (detailed in A) describes the task and clearly defines the four categories, providing one generic example for each. Additionally, the template specifies that the answer should be provided in a numeric code format to facilitate parsing and ensure unambiguous responses. The resulting prompts were then submitted to the LLM to obtain an answer for each pair of topics. To ensure a fair comparison, all models received the same prompt.
4.2 Chain-of-thought prompting
CoT prompting experiments operated similarly to the previous solution but involved a two-phase interaction with the LLM, as illustrated in Fig. 1(b). In the first phase, the model is asked to provide a definition of both topics, formulate a sentence that incorporates both of them, and finally discuss their potential semantic relationships. The resulting response is then fed back into the model in the second phase, with the sole task of determining the type of semantic relationship. The two templates used to produce the prompts for the chain-of-thought experiments are available in B.
4.3 One-way vs. two-way strategies
The experiments for both standard prompting and CoT prompting are conducted in two different configurations. The one-way strategy simply asks for the relationship between and . The two-way strategy involves asking the LLM to identify the relationship between topics and , as well as the relationship between topics and , when the order of the topics reversed. This straightforward approach helps to detect and potentially correct inconsistencies in the model’s responses. This is feasible because the broader and narrower are inverse relationships, while same-as and other are symmetric. For example, if the model determines that is broader than , but also that is same-as , we can infer that at least one of these answers is incorrect.
To manage these potential inconsistencies, we designed a set of empirical rules (cyan boxes as referees in both Fig. 1(a) and 1(b)). These rules were designed to prioritise the development of the hierarchical structure.
Considering f(X) and s(X) as the relationship types respectively obtained via the first and second branches of the two-way strategy; and len() as the length of the topic’s surface form, the rules are established as follows:
1.
broader :- f(broader) s(narrower)
2.
narrower :- f(narrower) s(broader)
3.
broader :- ((f(narrower) s(narrower)) (f(broader) s(broader))) len() len()
4.
narrower :- ((f(narrower) s(narrower)) (f(broader) s(broader))) len() len()
5.
same-as :- f(same-as) s(same-as)
6.
broader :- (f(broader) s(other)) (f(other) s(narrower))
7.
narrower :- (f(narrower) s(other)) (f(other) s(broader))
8.
:- f(X)
4.4 Experimental setup
To interact with the various LLMs we employed two services, e.g., Amazon Bedrock212121Amazon Bedrock - https://aws.amazon.com/bedrock, and OpenAI API222222OpenAI API - https://platform.openai.com/docs/overview; as well as the KoboldAI232323Kobold API - https://github.com/KoboldAI/KoboldAI-Client library for running LLMs locally.
Specifically, through Amazon Bedrock, we interacted with MistralAI family LLMs (e.g., Mistral-7B and Mixtral-87B), Anthropic models (e.g., Cluade 3 Haiku, and Claude 3 Sonnet), and Llama-2-70B by Meta. This facility provides a pre-structured generic wrapper, which encapsulates the interaction with the different LLMs offered by Amazon Bedrock.
With the OpenAI API we interacted with GPT models, such as GPT3.5-Turbo-Instruct and GPT4-Turbo.
Finally, we employed KoboldAI to interface with Dolphin-2.1-Mistral-7B, Mistral-7B-OpenOrca, Dolphin-2.6-Mistral-7B-dpo-laser, Dolphin-2.1-OpenOrca-7B, SOLAR-10.7B-Instruct-v1.0, OpenChat-3.5-1210, OpenChat-3.5-0106-Gemma, EURUS-7B-SFT. KoldbolAI is an open-source tool built on llama.cpp and provides model access through an API endpoint. We run it on a Google Colaboratory instance equipped with Nvidia’s V100 and L4 GPUs.
Within the GitHub repository, we report the various parameters we set when interacting with the models across the three services.
5 Results and Discussion
This section presents the results of the complete set of experiments described in Section 4. First we will discuss the standard prompting experiments (see Fig. 1(a) for flowchart) and then we will report the results of the chain-of-thought prompting experiments (see Fig. 1(b)). Further, we will compare the results across the various strategies.
5.1 Standard prompting, one-way strategy
Table 1 reports the result of the experiments with standard prompting when applying one-way strategy. Among the full open models (first three rows), llama-2 and mixtral demonstrate good performance in terms of average F1 with 0.669 and 0.779, respectively. In particular, mixtral is highly precise with the broader (0.819) and same-as (1), whereas llama-2 obtains high values of precision with narrower (1) and other (0.877). Complementarily, mixtral reaches high values of recall for narrower (0.888) and other (0.940), whereas llama-2 for broader (0.872) and same-as (0.992). In this case, mistral consistently performs low, except for the recall of the relationship other (0.992).
Among quantised models, dolphin-openorca and openchat stand out with the highest average F1-score of 0.724 and 0.705, respectively. While both consistently excel in precision, they face some challenges in recall for narrower and same-as relations. Other quantised models, although falling below 0.7 in average F1-score, still demonstrate remarkable performance in specific relationships. For instance, mistral-openorca excels in precision for the same-as (1) and other (0.992) relations and recall for narrower (0.996), despite lower performance in all other metrics.
When evaluating the proprietary models (as shown in the last four rows of Table 1), sonnet stands out, achieving an outstanding average F1-score of 0.967, with a well-balanced precision and recall across all relationship types. gpt-4 also demonstrates strong performance (0.952 F1). These two models exhibit high precision for same-as and other relations, and high recall for broader relations. In contrast, the two other models do not perform as well. Haiku, with an F1-score of 0.712, is surpassed by both mixtral and dolphin-openorca. Surprisingly, gpt-3.5 also shows lower-than-expected results.
Overall, we observe a significant performance gap between full open models and proprietary models. For example, sonnet outperforms mixtral, achieving an average F1-score that is 0.188 points higher.
| MODEL | PRECISION | RECALL | F1-SCORE | ||||||||||||
| AVG | BR | NR | SA | OT | AVG | BR | NR | SA | OT | AVG | BR | NR | SA | OT | |
| mistral | 0.401 | 0.667 | 0.000 | 0.511 | 0.428 | 0.462 | 0.008 | 0.000 | 0.848 | 0.992 | 0.313 | 0.016 | 0.000 | 0.638 | 0.598 |
| mixtral | 0.817 | 0.819 | 0.634 | 1.000 | 0.816 | 0.781 | 0.688 | 0.888 | 0.608 | 0.940 | 0.779 | 0.748 | 0.740 | 0.756 | 0.874 |
| llama-2 | 0.809 | 0.796 | 1.000 | 0.565 | 0.877 | 0.722 | 0.872 | 0.140 | 0.992 | 0.884 | 0.669 | 0.832 | 0.246 | 0.720 | 0.880 |
| dolphin-mistral | 0.676 | 0.613 | 0.736 | 0.901 | 0.455 | 0.599 | 0.652 | 0.412 | 0.508 | 0.824 | 0.599 | 0.632 | 0.528 | 0.650 | 0.586 |
| dolphin-mistral-dpo | 0.734 | 0.613 | 0.868 | 0.960 | 0.497 | 0.603 | 0.988 | 0.236 | 0.284 | 0.904 | 0.552 | 0.757 | 0.371 | 0.438 | 0.641 |
| dolphin-openorca | 0.780 | 0.612 | 0.810 | 0.960 | 0.737 | 0.733 | 0.952 | 0.496 | 0.576 | 0.908 | 0.724 | 0.745 | 0.615 | 0.720 | 0.814 |
| openchat | 0.826 | 0.930 | 0.991 | 0.902 | 0.481 | 0.701 | 0.740 | 0.416 | 0.664 | 0.984 | 0.705 | 0.824 | 0.586 | 0.765 | 0.646 |
| openchat-gemma | 0.679 | 0.396 | 0.471 | 1.000 | 0.849 | 0.523 | 0.968 | 0.296 | 0.268 | 0.560 | 0.506 | 0.562 | 0.364 | 0.423 | 0.675 |
| solar | 0.723 | 0.642 | 0.877 | 0.850 | 0.523 | 0.657 | 0.916 | 0.544 | 0.784 | 0.384 | 0.646 | 0.755 | 0.672 | 0.529 | 0.627 |
| mistral-openorca | 0.815 | 0.866 | 0.401 | 1.000 | 0.992 | 0.605 | 0.568 | 0.996 | 0.384 | 0.472 | 0.613 | 0.686 | 0.572 | 0.555 | 0.640 |
| eurus | 0.596 | 0.410 | 0.638 | 0.424 | 0.911 | 0.493 | 0.984 | 0.204 | 0.332 | 0.452 | 0.466 | 0.579 | 0.309 | 0.372 | 0.604 |
| llama-3 | 0.715 | 0.560 | 0.446 | 1.000 | 0.853 | 0.582 | 0.732 | 0.808 | 0.256 | 0.532 | 0.568 | 0.634 | 0.575 | 0.408 | 0.655 |
| orca-2 | 0.500 | 0.342 | 0.382 | 0.486 | 0.788 | 0.456 | 0.944 | 0.052 | 0.068 | 0.760 | 0.372 | 0.502 | 0.091 | 0.119 | 0.774 |
| gpt-3.5 | 0.668 | 0.267 | 0.407 | 1.000 | 1.000 | 0.295 | 0.976 | 0.096 | 0.032 | 0.076 | 0.195 | 0.419 | 0.155 | 0.062 | 0.141 |
| gpt-4 | 0.952 | 0.919 | 0.934 | 1.000 | 0.954 | 0.949 | 1.000 | 0.968 | 0.832 | 0.996 | 0.948 | 0.958 | 0.951 | 0.908 | 0.975 |
| haiku | 0.820 | 0.496 | 0.870 | 1.000 | 0.912 | 0.710 | 0.980 | 0.564 | 0.468 | 0.828 | 0.712 | 0.659 | 0.684 | 0.638 | 0.868 |
| sonnet | 0.969 | 0.929 | 0.950 | 0.996 | 1.000 | 0.967 | 0.996 | 0.996 | 0.932 | 0.944 | 0.967 | 0.961 | 0.973 | 0.963 | 0.971 |
5.2 Standard prompting, two-way strategy
Table 2 presents the performance of the models when using standard prompting with the two-way strategy.
Among the full open models, mixtral achieves an average F1-score of 0.847, significantly outperforming both llama-2 and mistral. It is able to identify the broader, narrower, and same-as with F1-score around 0.82, but shows exceptional capabilities (0.925) in determining the relationship other.
Regarding the quantised models, there are three models (e.g., dolphin-mistral-dpo, dolphin-openorca, and openchat) that obtain high values of average F1-score above 0.78.
Among the proprietary models, sonnet obtained comparable results to previous experiments (0.965 F1), where the performance of gpt-3.5, gpt-4, and haiku obtained a slight improvement.
In this case, the full open models and their quantised counterparts exhibit similar performance levels (e.g., 0.847 for mixtral and 0.853 for dolphin-openorca). Moreover, the two-way strategy has effectively reduced the performance gap between these models and their larger, proprietary counterparts, narrowing the difference to just 0.113 F1 points.
| MODEL | PRECISION | RECALL | F1-SCORE | ||||||||||||
| AVG | BR | NR | SA | OT | AVG | BR | NR | SA | OT | AVG | BR | NR | SA | OT | |
| mistral | 0.239 | 0.000 | 0.000 | 0.451 | 0.503 | 0.475 | 0.000 | 0.000 | 0.908 | 0.992 | 0.318 | 0.000 | 0.000 | 0.603 | 0.668 |
| mixtral | 0.865 | 0.751 | 0.773 | 1.000 | 0.935 | 0.846 | 0.904 | 0.856 | 0.708 | 0.916 | 0.847 | 0.820 | 0.812 | 0.829 | 0.925 |
| llama-2 | 0.759 | 0.832 | 0.819 | 0.426 | 0.959 | 0.621 | 0.336 | 0.308 | 0.996 | 0.844 | 0.605 | 0.479 | 0.448 | 0.596 | 0.898 |
| dolphin-mistral | 0.718 | 0.686 | 0.673 | 0.884 | 0.629 | 0.704 | 0.708 | 0.732 | 0.704 | 0.672 | 0.706 | 0.697 | 0.701 | 0.764 | 0.664 |
| dolphin-mistral-dpo | 0.834 | 0.759 | 0.764 | 0.968 | 0.844 | 0.808 | 0.956 | 0.932 | 0.476 | 0.868 | 0.795 | 0.846 | 0.840 | 0.638 | 0.856 |
| dolphin-openorca | 0.862 | 0.799 | 0.774 | 0.928 | 0.947 | 0.852 | 0.888 | 0.892 | 0.776 | 0.852 | 0.853 | 0.841 | 0.829 | 0.845 | 0.897 |
| openchat | 0.838 | 0.979 | 0.964 | 0.823 | 0.586 | 0.777 | 0.736 | 0.648 | 0.744 | 0.980 | 0.783 | 0.840 | 0.775 | 0.782 | 0.734 |
| openchat-gemma | 0.734 | 0.484 | 0.512 | 1.000 | 0.939 | 0.590 | 0.804 | 0.792 | 0.332 | 0.432 | 0.579 | 0.605 | 0.622 | 0.499 | 0.592 |
| solar | 0.723 | 0.731 | 0.753 | 0.784 | 0.623 | 0.715 | 0.836 | 0.816 | 0.508 | 0.700 | 0.710 | 0.780 | 0.783 | 0.617 | 0.659 |
| mistral-openorca | 0.781 | 0.542 | 0.584 | 1.000 | 1.000 | 0.650 | 0.888 | 0.908 | 0.492 | 0.312 | 0.630 | 0.673 | 0.711 | 0.660 | 0.476 |
| eurus | 0.587 | 0.480 | 0.495 | 0.374 | 1.000 | 0.489 | 0.584 | 0.576 | 0.492 | 0.304 | 0.488 | 0.527 | 0.532 | 0.425 | 0.466 |
| llama-3 | 0.746 | 0.476 | 0.544 | 1.000 | 0.964 | 0.586 | 0.856 | 0.836 | 0.328 | 0.324 | 0.562 | 0.611 | 0.659 | 0.494 | 0.485 |
| orca-2 | 0.561 | 0.429 | 0.440 | 0.434 | 0.940 | 0.519 | 0.668 | 0.688 | 0.092 | 0.628 | 0.491 | 0.523 | 0.537 | 0.152 | 0.753 |
| gpt-3.5 | 0.690 | 0.360 | 0.401 | 1.000 | 1.000 | 0.389 | 0.748 | 0.744 | 0.032 | 0.032 | 0.283 | 0.486 | 0.521 | 0.062 | 0.062 |
| gpt-4 | 0.965 | 0.936 | 0.936 | 1.000 | 0.988 | 0.963 | 1.000 | 0.996 | 0.868 | 0.988 | 0.962 | 0.967 | 0.965 | 0.929 | 0.988 |
| haiku | 0.842 | 0.704 | 0.668 | 0.994 | 1.000 | 0.792 | 0.896 | 0.908 | 0.640 | 0.724 | 0.794 | 0.789 | 0.769 | 0.779 | 0.840 |
| sonnet | 0.967 | 0.929 | 0.940 | 1.000 | 1.000 | 0.965 | 0.992 | 0.996 | 0.948 | 0.924 | 0.965 | 0.959 | 0.967 | 0.973 | 0.961 |
5.3 Chain-of-thought prompting, one-way strategy
Table 3 presents the results of the chain-of-thought, one-way strategy experiments. All models demonstrate significantly better performance compared to the analogous experiment using standard prompting. This confirms that CoT is a highly effective technique for enhancing model efficacy. A notable phenomenon observed here is that almost all models exhibit high precision on the same-as relationship. However, this often comes at the cost of recall, with the highest recall achieved by openchat, reaching 0.780.
Among full open models, mixtral again outperformed the other model with an average F1-score of 0.808.
Regarding the quantised models, dolphin-mistral achieves the highest average F1-score of 0.869. It is closely followed by dolphin-mistral-dpo with a score of 0.826, solar at 0.825, and openchat at 0.823. This suggests that models fine-tuned on the OpenOrca corpus are particularly effective for this task. In this context, the use of the CoT strategy enables the quantised model dolphin-mistral to even surpass the much larger mixtral model.
Among the proprietary models, gpt-4 achieved the highest performance, reaching an average F1-score of 0.911. However, both gpt-4 and sonnet show a slight decrease in performance compared to other models. This suggests that some of the more advanced models might be confused rather than aided by the CoT strategy. On the other hand, both haiku and gpt-3.5 perform significantly better. Of course, without detailed knowledge of the pipelines of these more modern proprietary models, and whether they benefit from additional systems, it is difficult to draw definitive conclusions.
| MODEL | PRECISION | RECALL | F1-SCORE | ||||||||||||
| AVG | BR | NR | SA | OT | AVG | BR | NR | SA | OT | AVG | BR | NR | SA | OT | |
| mistral | 0.766 | 0.667 | 0.749 | 0.956 | 0.695 | 0.720 | 0.936 | 0.632 | 0.348 | 0.964 | 0.696 | 0.779 | 0.685 | 0.510 | 0.807 |
| mixtral | 0.849 | 0.732 | 0.807 | 1.000 | 0.858 | 0.820 | 0.948 | 0.904 | 0.488 | 0.940 | 0.808 | 0.826 | 0.853 | 0.656 | 0.897 |
| llama-2 | 0.775 | 0.466 | 0.674 | 1.000 | 0.960 | 0.612 | 0.976 | 0.860 | 0.224 | 0.388 | 0.576 | 0.630 | 0.756 | 0.366 | 0.553 |
| dolphin-mistral | 0.892 | 0.749 | 0.912 | 1.000 | 0.907 | 0.871 | 1.000 | 0.872 | 0.680 | 0.932 | 0.869 | 0.856 | 0.892 | 0.810 | 0.919 |
| dolphin-mistral-dpo | 0.863 | 0.685 | 0.947 | 0.994 | 0.827 | 0.827 | 0.992 | 0.708 | 0.688 | 0.920 | 0.826 | 0.811 | 0.810 | 0.813 | 0.871 |
| dolphin-openorca | 0.840 | 0.524 | 0.865 | 1.000 | 0.971 | 0.736 | 1.000 | 0.848 | 0.560 | 0.536 | 0.738 | 0.688 | 0.857 | 0.718 | 0.691 |
| openchat | 0.844 | 0.692 | 0.874 | 0.947 | 0.864 | 0.822 | 0.988 | 0.780 | 0.780 | 0.740 | 0.823 | 0.814 | 0.825 | 0.855 | 0.797 |
| openchat-gemma | 0.822 | 0.429 | 0.864 | 1.000 | 0.993 | 0.650 | 1.000 | 0.408 | 0.588 | 0.604 | 0.662 | 0.600 | 0.554 | 0.741 | 0.751 |
| solar | 0.858 | 0.739 | 0.771 | 1.000 | 0.924 | 0.830 | 0.940 | 0.928 | 0.580 | 0.872 | 0.825 | 0.828 | 0.842 | 0.734 | 0.897 |
| mistral-openorca | 0.810 | 0.551 | 0.865 | 0.987 | 0.836 | 0.736 | 1.000 | 0.744 | 0.608 | 0.592 | 0.739 | 0.710 | 0.800 | 0.753 | 0.693 |
| eurus | 0.796 | 0.661 | 0.635 | 1.000 | 0.887 | 0.723 | 1.000 | 0.912 | 0.416 | 0.564 | 0.706 | 0.796 | 0.749 | 0.588 | 0.690 |
| llama-3 | 0.723 | 0.547 | 0.491 | 1.000 | 0.852 | 0.626 | 0.696 | 0.748 | 0.232 | 0.828 | 0.605 | 0.613 | 0.593 | 0.377 | 0.840 |
| orca-2 | 0.787 | 0.622 | 0.624 | 1.000 | 0.904 | 0.711 | 0.868 | 0.928 | 0.404 | 0.644 | 0.700 | 0.725 | 0.746 | 0.576 | 0.752 |
| gpt-3.5 | 0.727 | 0.534 | 0.501 | 0.896 | 0.976 | 0.573 | 0.904 | 0.880 | 0.344 | 0.164 | 0.522 | 0.672 | 0.639 | 0.497 | 0.281 |
| gpt-4 | 0.924 | 0.828 | 0.921 | 1.000 | 0.945 | 0.914 | 0.984 | 0.980 | 0.724 | 0.968 | 0.911 | 0.899 | 0.950 | 0.840 | 0.957 |
| haiku | 0.861 | 0.664 | 0.809 | 1.000 | 0.969 | 0.812 | 0.996 | 0.952 | 0.540 | 0.760 | 0.806 | 0.797 | 0.875 | 0.701 | 0.852 |
| sonnet | 0.909 | 0.821 | 0.865 | 1.000 | 0.950 | 0.896 | 0.992 | 0.976 | 0.696 | 0.920 | 0.893 | 0.899 | 0.917 | 0.821 | 0.935 |
5.4 Chain-of-thought prompting, two-way strategy
The values are generally consistent with those obtained from the chain-of-thought one-way strategy experiments (see Table 3). Specifically, mixtral, dolphin-mistral, and gtp-4 emerge as the top-performing models among the full open models, the quantised models, and the proprietary models, respectively. The average F1 score increases by a few percentage points in 15 out of 17 models, confirming that the two-way strategy remains effective when combined with CoT.
| MODEL | PRECISION | RECALL | F1-SCORE | ||||||||||||
| AVG | BR | NR | SA | OT | AVG | BR | NR | SA | OT | AVG | BR | NR | SA | OT | |
| mistral | 0.830 | 0.775 | 0.734 | 0.932 | 0.880 | 0.810 | 0.936 | 0.904 | 0.492 | 0.908 | 0.799 | 0.848 | 0.810 | 0.644 | 0.894 |
| mixtral | 0.876 | 0.768 | 0.795 | 1.000 | 0.940 | 0.850 | 0.992 | 0.948 | 0.576 | 0.884 | 0.843 | 0.866 | 0.865 | 0.731 | 0.911 |
| llama-2 | 0.766 | 0.502 | 0.563 | 1.000 | 1.000 | 0.588 | 0.936 | 0.928 | 0.280 | 0.208 | 0.534 | 0.654 | 0.701 | 0.438 | 0.344 |
| dolphin-mistral | 0.930 | 0.874 | 0.852 | 1.000 | 0.995 | 0.920 | 0.996 | 0.992 | 0.844 | 0.848 | 0.920 | 0.931 | 0.917 | 0.915 | 0.916 |
| dolphin-mistral-dpo | 0.915 | 0.859 | 0.831 | 0.986 | 0.982 | 0.905 | 0.952 | 0.964 | 0.840 | 0.864 | 0.906 | 0.903 | 0.893 | 0.907 | 0.919 |
| dolphin-openorca | 0.828 | 0.640 | 0.670 | 1.000 | 1.000 | 0.745 | 0.968 | 0.968 | 0.704 | 0.340 | 0.724 | 0.771 | 0.792 | 0.826 | 0.508 |
| openchat | 0.857 | 0.763 | 0.777 | 0.903 | 0.986 | 0.834 | 0.940 | 0.932 | 0.892 | 0.572 | 0.828 | 0.842 | 0.847 | 0.897 | 0.724 |
| openchat-gemma | 0.798 | 0.601 | 0.597 | 0.994 | 1.000 | 0.711 | 0.868 | 0.852 | 0.672 | 0.452 | 0.709 | 0.710 | 0.702 | 0.802 | 0.623 |
| solar | 0.886 | 0.759 | 0.815 | 1.000 | 0.971 | 0.861 | 0.980 | 0.968 | 0.704 | 0.792 | 0.860 | 0.855 | 0.885 | 0.826 | 0.872 |
| mistral-openorca | 0.846 | 0.686 | 0.698 | 1.000 | 1.000 | 0.784 | 0.968 | 0.972 | 0.756 | 0.440 | 0.772 | 0.803 | 0.813 | 0.861 | 0.611 |
| eurus | 0.833 | 0.672 | 0.670 | 1.000 | 0.990 | 0.756 | 0.992 | 0.992 | 0.628 | 0.412 | 0.739 | 0.801 | 0.800 | 0.772 | 0.582 |
| llama-3 | 0.766 | 0.556 | 0.541 | 1.000 | 0.968 | 0.661 | 0.808 | 0.812 | 0.292 | 0.732 | 0.649 | 0.659 | 0.650 | 0.452 | 0.834 |
| orca-2 | 0.809 | 0.598 | 0.651 | 1.000 | 0.985 | 0.724 | 0.904 | 0.912 | 0.544 | 0.536 | 0.720 | 0.720 | 0.760 | 0.705 | 0.694 |
| gpt-3.5 | 0.741 | 0.547 | 0.553 | 0.863 | 1.000 | 0.606 | 0.904 | 0.920 | 0.528 | 0.072 | 0.540 | 0.682 | 0.691 | 0.655 | 0.134 |
| gpt-4 | 0.941 | 0.899 | 0.889 | 1.000 | 0.975 | 0.935 | 1.000 | 0.996 | 0.792 | 0.952 | 0.933 | 0.947 | 0.940 | 0.884 | 0.964 |
| haiku | 0.877 | 0.755 | 0.765 | 1.000 | 0.989 | 0.842 | 0.984 | 0.988 | 0.680 | 0.716 | 0.839 | 0.854 | 0.862 | 0.809 | 0.831 |
| sonnet | 0.916 | 0.822 | 0.855 | 1.000 | 0.986 | 0.901 | 1.000 | 0.992 | 0.768 | 0.844 | 0.900 | 0.902 | 0.918 | 0.869 | 0.909 |
5.5 Comparative Analysis
Table 5 presents the average F1-scores for all 17 models across the four strategies, including the relative differences between one-way and two-way strategies for both standard and chain-of-thought prompting. The results indicate that shifting from a one-way to a two-way strategy improved performance in 14 models for standard prompting and in 15 models for chain-of-thought prompting.
Generally, we observe a gradual increase in the average F1-score from left to right in Table 5, demonstrating the beneficial effects of both CoT and the two-way strategy. Thirteen models showed overall improvement, while four models (llama-2, dolphin-openorca, gpt-4, and sonnet) experienced a decline in performance. When applying a 0.7 average F1-score as a quality benchmark, only dolphin-openorca and openchat surpassed this threshold under a standard prompting approach with a one-way strategy. However, with a two-way strategy in standard prompting, five models exceeded the threshold. This number increased to eight models when a chain-of-thought, one-way strategy was used. Ultimately, nine models achieved an average F1-score above the 0.7 threshold when employing a chain-of-thought, two-way strategy. These results underscore the critical importance of a good prompting strategy when working in a zero-shot setting.
Finally, we can analyse the best model for each of the three categories: fully open models, quantised models, and proprietary models. The mixtral model stands out among fully open models, particularly when used with the standard prompting and two-way strategy (0.847 F1). In contrast, the dolphin-mistral model surpasses other quantised models when applied with the chain-of-thought prompting and two-way strategy (0.920 F1). Notably, dolphin-mistral outperforms mixtral, likely due to the additional fine-tuning on datasets like OpenOrca, which seem particularly beneficial for this task. Among proprietary models, sonnet delivers the best performance when using standard prompts and a straightforward, one-way approach (0.967 F1). Interestingly, unlike most open models, its performance declines with more complex prompts. Since sonnet is a black-box system that may incorporate additional components, pinpointing the exact cause of this behaviour is challenging.
| MODEL | Standard Prompting | CoT Prompting | ||||
| One-way | Two-way | Difference | One-way | Two-way | Difference | |
| mistral | 0.313 | 0.318 | 0.005 | 0.696 | 0.799 | 0.103 |
| mixtral | 0.779 | 0.847 | 0.068 | 0.808 | 0.843 | 0.035 |
| llama-2 | 0.669 | 0.605 | -0.064 | 0.576 | 0.534 | -0.042 |
| dolphin-mistral | 0.599 | 0.706 | 0.107 | 0.869 | 0.920 | 0.051 |
| dolphin-mistral-dpo | 0.552 | 0.795 | 0.243 | 0.826 | 0.906 | 0.080 |
| dolphin-openorca | 0.724 | 0.853 | 0.129 | 0.738 | 0.724 | -0.014 |
| openchat | 0.705 | 0.783 | 0.078 | 0.823 | 0.828 | 0.005 |
| openchat-gemma | 0.506 | 0.579 | 0.073 | 0.662 | 0.709 | 0.047 |
| solar | 0.646 | 0.710 | 0.064 | 0.825 | 0.860 | 0.035 |
| mistral-openorca | 0.613 | 0.630 | 0.017 | 0.739 | 0.772 | 0.033 |
| eurus | 0.466 | 0.488 | 0.022 | 0.706 | 0.739 | 0.033 |
| llama-3 | 0.568 | 0.562 | -0.006 | 0.605 | 0.649 | 0.044 |
| orca-2 | 0.372 | 0.491 | 0.119 | 0.700 | 0.720 | 0.020 |
| gpt-3.5 | 0.195 | 0.283 | 0.088 | 0.522 | 0.540 | 0.018 |
| gpt-4 | 0.948 | 0.962 | 0.014 | 0.911 | 0.933 | 0.022 |
| haiku | 0.712 | 0.794 | 0.082 | 0.806 | 0.839 | 0.033 |
| sonnet | 0.967 | 0.965 | -0.002 | 0.893 | 0.900 | 0.007 |
6 Error Analysis
In this section, we examine common error types displayed by the top-performing LLMs across the three categories: a) mixtral (standard prompting, two-way), b) dolphin-mistral (CoT, two-way), and c) sonnet (standard prompting, one-way). Figure 2 reports the relevant confusion matrices.
two-way strategy.
two-way strategy.
one-way strategy.
The mixtral model misclassified 154 out of 1,000 topic pairs. Specifically, 75 pairs were incorrectly identified as broader, 63 as narrower, 16 as other, and none as same-as. mixtral demonstrates a highly conservative approach in predicting the same-as category, achieving a precision of 1, as indicated in Table 2. This comes at a cost, as the model exhibits a tendency to misclassify pairs that are labelled as same-as in the gold standard, often categorising them incorrectly as either broader or narrower. This misclassification results in a low recall for the same-as category (0.708), which represents a significant limitation of the model. Specifically, 29.2% of same-as pairs were misclassified: 16% were labelled as broader, 12.4% as narrower, and 0.8% as other. Additionally, the mixtral model occasionally confuses broader pairs with narrower ones and vice versa. Notably, 9.6% of broader pairs were incorrectly predicted, with 8% labelled as narrower and 1.6% as other. Similarly, 14.4% of narrower pairs were misclassified, with 10.4% labelled as broader and 4% as other.
The dolphin-mistral model exhibits a different pattern compared to mixtral, particularly excelling in pairs labelled as broader and narrower in the gold standard. It makes 80 misclassifications, predominantly in the same-as (39) and other (38) categories. Although the model still faces challenges with same-as relationships, it performs better than mixtral. Specifically, for the same-as relationships, a small percentage (6.8%) are misclassified as broader, while 8.8% are incorrectly identified as narrower.
The sonnet model exhibits a similar performance pattern to dolphin-mistral, achieving near-perfect accuracy across both broader and narrow categories, while showing notable improvements in the other two categories. The model made only 33 classification errors, with the majority occurring in same-as (17) and other (14). Specifically, within the same-as misclassifications, 6% were incorrectly identified as broader, and 0.8% as narrower. In contrast, for the other category, 1.6% of errors were misclassified as broader or same-as, while 4% were erroneously categorised as other.
The error analysis provides several valuable insights. First, the best-performing model excels in identifying hierarchical relationships, such as broader and narrower. This is a promising indication of the potential for LLMs to assist in ontology generation, where hierarchical relationships are fundamental to structuring the ontology. Second, while the very best model shows reasonable success in identifying “same-as” relationships, this task proves challenging. This difficulty may arise because same-as relationships between research topics are not purely syntactic synonyms and need deep expertise in the field to be interpreted correctly. Fine-tuning the model on task-specific data could help address this issue, and we plan to explore this approach in future work.
7 Conclusions
This paper conducted an in-depth analysis of the capability of LLMs to identify semantic relationships between research topics for the purpose of ontology generation. We evaluated seventeen LLMs, categorised into three groups: fully open models (3), open quantised models (10), and proprietary models (4). We evaluated their performance against a gold standard of 1,000 relationships from the IEEE Thesaurus. We also investigated four distinct zero-shot prompting strategies, which yielded significantly different outcomes.
Our results demonstrated the impressive zero-shot reasoning capabilities of state-of-the-art models. Three models emerged as top performers within their respective categories: Mixtral-8×7B (mixtral: 0.847 average F1-score) for the full open models; Dolphin-2.1-Mistral-7B (dolphin-mistral: 0.920) for open quantised models; and Claude 3 Sonnet (sonnet: 0.967) for proprietary models.
We are currently working on several fronts. We are fine-tuning off-the-shelf models like BERT, Mistral, LLaMa 3, Flan T5, Gemma, and Phi 3 with the aim of enhancing the performance of lightweight models to reduce computational costs and environmental impact when generating ontologies of research areas. Furthermore, we plan to incorporate a semantic reasoner, such as RDFox, into our pipeline for identifying and resolving inconsistencies in the generated ontology. Finally, we aim to extend our analysis to other scientific fields such as Material Science, Medicine, and Physics, with the ultimate goal of creating a comprehensive, multi-field ontology of research topics.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used Grammarly in order to improve readability. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
Appendix A Prompt for Standard Prompting
For consistency we applied the same prompt across all the models. We engineered this prompt through various refinements, to ensure optimal comprehension of the task and accurate responses. Below is the template of our prompt, customised for each topic pair by substituting [TOPIC-A] for the first topic and [TOPIC-B] for the second.
Classify the relationship between ’[TOPIC-A]’ and ’[TOPIC-B]’ by applying the following relationship definitions:1. ’[TOPIC-A]’ is-broader-than ’[TOPIC-B]’ if ’[TOPIC-A]’ is a super-category of ’[TOPIC-B]’, that is ’[TOPIC-B]’ is a type, a branch, or a specialised aspect of ’[TOPIC-A]’ or that ’[TOPIC-B]’ is a tool or a methodology mostly used in the context of ’[TOPIC-A]’ (e.g., car is-broader-than wheel).2. ’[TOPIC-A]’ is-narrower-than ’[TOPIC-B]’ if ’[TOPIC-A]’ is a sub-category of ’[TOPIC-B]’, that is ’[TOPIC-A]’ is a type, a branch, or a specialised aspect of ’[TOPIC-B]’ or that ’[TOPIC-A]’ is a tool or a methodology mostly used in the context of ’[TOPIC-B]’ (e.g., wheel is-narrower-than car).3. ’[TOPIC-A]’ is-same-as-than ’[TOPIC-B]’ if ’[TOPIC-A]’ and ’[TOPIC-B]’ are synonymous terms denoting an identical concept (e.g., beautiful is-same-as-than attractive), including when one is the plural form of the other (e.g., cat is-same-as-than cats).4. ’[TOPIC-A]’ is-other-than ’[TOPIC-B]’ if ’[TOPIC-A]’ and ’[TOPIC-B]’ either have no direct relationship or share a different kind of relationship that does not fit into the other defined relationships.Given the previous definitions, determine which one of the following statements is correct:1. ’[TOPIC-A]’ is-broader-than ’[TOPIC-B]’2. ’[TOPIC-B]’ is-narrower-than ’[TOPIC-A]’3. ’[TOPIC-A]’ is-narrower-than ’[TOPIC-B]’4. ’[TOPIC-B]’ is-broader-than ’[TOPIC-A]’5. ’[TOPIC-A]’ is-same-as-than ’[TOPIC-B]’6. ’[TOPIC-A]’ is-other-than ’[TOPIC-B]’Answer by only stating the correct statement and its number.
Appendix B Prompt for Chain-of-Thought Prompting
In the CoT prompting we have a two-phase interaction with the model. In the first interaction the model is sked to provide a definition of both topics, formulate a sentence that incorporates both topics, and finally discuss their potential semantic relationships. The resulting response is then fed back into the model in the second phase, with the sole task of determining the type of semantic relationship. Below are our prompt templates, customised for each topic pair by substituting [TOPIC-A] for the first topic and [TOPIC-B] for the second.
First interaction
Classify the relationship between ’[TOPIC-A]’ and ’[TOPIC-B]’ by applying the following relationship definitions:1. ’[TOPIC-A]’ is-broader-than ’[TOPIC-B]’ if ’[TOPIC-A]’ is a super-category of ’[TOPIC-B]’, that is ’[TOPIC-B]’ is a type, a branch, or a specialised aspect of ’[TOPIC-A]’ or that ’[TOPIC-B]’ is a tool or a methodology mostly used in the context of ’[TOPIC-A]’ (e.g., car is-broader-than wheel).2. ’[TOPIC-A]’ is-narrower-than ’[TOPIC-B]’ if ’[TOPIC-A]’ is a sub-category of ’[TOPIC-B]’, that is ’[TOPIC-A]’ is a type, a branch, or a specialised aspect of ’[TOPIC-B]’ or that ’[TOPIC-A]’ is a tool or a methodology mostly used in the context of ’[TOPIC-B]’ (e.g., wheel is-narrower-than car).3. ’[TOPIC-A]’ is-same-as-than ’[TOPIC-B]’ if ’[TOPIC-A]’ and ’[TOPIC-B]’ are synonymous terms denoting a very similar concept (e.g., ’beautiful’ is-same-as-than ’attractive’), including when one is the plural form of the other (e.g., cat is-same-as-than cats).4. ’[TOPIC-A]’ is-other-than ’[TOPIC-B]’ if ’[TOPIC-A]’ and ’[TOPIC-B]’ either have no direct relationship or share a different kind of relationship that does not fit into the other defined relationships.Think step by step by following these sequential instructions:1) Provide a precise definition for ’[TOPIC-A]’.2) Provide a precise definition for ’[TOPIC-B]’.3) Formulate a sentence that includes both ’[TOPIC-A]’ and ’[TOPIC-B]’.4) Discuss ’[TOPIC-A]’ and ’[TOPIC-B]’ usage and relationship (is-narrower-than, is-broader-than, is-same-as-than, or is-other-than).
Second interaction
[PREVIOUS-RESPONSE]Given the previous discussion, determine which one of the following statements is correct:1. ’[TOPIC-A]’ is-broader-than ’[TOPIC-B]’2. ’[TOPIC-B]’ is-narrower-than ’[TOPIC-A]’3. ’[TOPIC-A]’ is-narrower-than ’[TOPIC-B]’4. ’[TOPIC-B]’ is-broader-than ’[TOPIC-A]’5. ’[TOPIC-A]’ is-same-as-than ’[TOPIC-B]’6. ’[TOPIC-A]’ is-other-than ’[TOPIC-B]’Answer by only stating the number of the correct statement.